OpenStudy (anonymous):

Use method of residues to find the partial fraction decomposition of e^-s+s/s^2+s+1

OpenStudy (anonymous):

Use method of residues to find the partial fraction decomposition of \[\frac{e^{-s}+s}{s^2+s+1}\]

OpenStudy (anonymous):

not sure how residues does it, but you can go head and use complex numbers to do it

OpenStudy (anonymous):

My main problem here is determining how to handle the irreducible quadratic with this method. I know if it was just something like s^2+1 we could split it as (s+i)(s-i). But I keep getting a different equation when I redistribute the factors I try

OpenStudy (anonymous):

you can solve using the quadratic formula, or recognize that this is what you get when you solve \[z^3=1\] i.e. \[z^3-1=0\]

OpenStudy (anonymous):

solve \[z^3-1=0\] gives \[(z-1)(z^2+z+1)=0\] the three roots of unity are \[\large \{1,e^{\frac{2\pi}{3}i},e^{\frac{4\pi}{3}i}\}\]

OpenStudy (anonymous):

probably easier to use exponential form than rectangular form, but if you use the quadratic formula you get \[\frac{-1\pm\sqrt3 i}{2}\] as your zeros same thing

OpenStudy (anonymous):

Okay. I think I understand most of that. Thanks!

OpenStudy (anonymous):

yw if the roots of unity business messed you up, then the quadratic formula always worked i just recognized that \(z^3-1\) factors as \((z-1)(z^2+z+1)\) so knew what the other two roots are quickly good luck